Exercise 8-7 Ceradyne Tax Service prepares tax returns for small businesses. The cost of preparing 800 tax returns in the prior year was: Direct labor $ 332,000 Variable overhead 312,000 Fixed overhead 248,000 Total cost $ 892,000 At the start of the current year, the company received an offer from Advantage Business, a firm that provides bundled services to businesses. Advantage wants Ceradyne Tax Service to prepare tax returns for its 120 small-business clients. Ceradyne Tax Service has the capacity to prepare up to 1,000 returns in a given year, so this special order would not take away revenue from any of Ceradyne Tax Service’s current clients. Advantage is willing to pay $900 per tax return. What will be the effect on Ceradyne Tax Service’s profit if it agrees to prepare returns for the 120 clients of Advantage Business? (Enter decrease in profit using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).) Profit will by $ per year.
Answer #1
Incremental revenue 108000 =120*900
Less: Incremental costs
Direct labor 49800 =332000/800*120
Variable overhead 46800 =312000/800*120
Total Incremental costs 96600
Incremental income 11400
Profit will increase by $11400 per year.